# Extreme events at the onset of epileptic-like chimeras in small-world networks of FitzHugh-Nagumo neurons

### v1

@@ II.1 Emergence of Epileptic-like events
In the following, we study more generally the region of existence of chimeras, and in particular, the transitions between the distinct dynamic regimes.
Figure 3: Emergence of epileptic-like states. (a)-(c) Frequency of epileptic-like events in $(N,d)$-parameter space, for (a) $p=0.1$, (b) $p=0.3$, and (c) $p=1.0$. (d) Shows the distribution of durations of the epileptic-like events in log-linear and (inset) log-log plots, and (e) shows the distribution of waiting times between epileptic-like events. Time corresponds to arbitrary simulation time units. Circles in panel (c) correspond to the system parameters used for (d) and (e).
II.2 Characterising the N-d-p phase space
@@
As an initial measure of the collective behaviour of the network of FHN coupled oscillators, we compute the time-averaged global order parameter ($\bar{R}$) as a function of the coupling strength, in a WS networks of different sizes and $p=1$ (see Fig. 4a). Here we note that regardless of the network size, $\bar{R}$ grows monotonically with $d$, exhibiting two transitions: One at $d\sim 0.05$ and another at $d\sim 0.15$, after which the system reaches full synchronisation. Notably, in the regions where chimeras are expected (including epileptic-like events) i.e. $d\sim[0.05,0.15]$), the averaged order parameter exhibits a plateau for large-enough system sizes.
Decreasing the rewiring probability, $p$, of the Watts-Strogatz network causes a significant increase in the coupling threshold at which the oscillators become fully synchronized (see Fig. 4b-c). This was consistent with an increase in the region of observation of epileptic-like events (see Figs. 3b-c), however this comes with a reduction in their frequency of appearance. To understand the changes in the transition line to synchronisation we performed a linear stability analysis of the synchronized solution.
@@
Figure 4: Phase transitions of coupled FHN oscillators. (a) Time-averaged order parameter, $\bar{R}$, as a function of the coupling strength, $d$, for WS networks with rewiring probability $p=1$ and varied sizes, $N$. (b)-(d) $\bar{R}$ in the $(N,d)$-parameter space, for (b) $p=0.1$, (c) $p=0.3$, and (d) $p=1.0$.
@@ II.2.1 Linear stability analysis of the synchronous solution
A linear stability analysis for the synchronised solution can be achieved by evaluating the Master Stability Function (MSF) of Eq. (1) [34, 3, 35] (see derivation in Appendix B). The MSF is defined as the maximum Lyapunov exponent $\Lambda_{max}(\nu)$ associated to the linearised dynamical system describing the time evolution of the perturbations transverse to the synchronous manifold $\mathcal{S}$, and allow us to study the stability of $\mathcal{S}$ as function of a single parameter $\nu=\gamma_{2}d$, where $d$ is the coupling strength and $\gamma_{2}$ is the second smallest eigenvalue of the network Laplacian matrix, known in the literature as algebraic connectivity [36]. When $\Lambda_{max}(\nu)>0$, $\mathcal{S}$ is unstable, while $\Lambda_{max}(\nu)<0$ provides a necessary condition for the stability of $\mathcal{S}$ [3]. In Fig. 5a, we show $\Lambda_{max}(\nu)$ for $\alpha=\pi/2-0.1$, which exhibits a positive window for the MSF, and for $\alpha=0$, for which the MSF is always negative and neither chimeras nor incoherent behaviour can be observed.
Consistent with the observations of the time-averaged global order parameter (Fig. 4b-d), lowering the rewiring probability $p$ causes a shift in the critical curve $\Lambda_{max}(\nu)=0$ towards larger values of $d$ (see Fig. 5b-d). We note that the critical curve was accurately predicted by the Wu-Chua conjecture (see Appendix C), delimiting the region above (below) which the synchronised manifold is stable (unstable), see dashed line in 5b-d.
The MSF analysis shows that the change in the transition line to synchronisation comes from changes in the topological features of the network, which enter in the MSF through the second smallest eigenvalue of the Laplacian matrix of the connectivity network. With this, we have been able to characterize the transition from intermittency (where chimeras are observed) to synchronisation, corresponding to the stabilization of the two-dimensional synchronisation manifold.
Figure 5: MSF of FHN coupled oscillators. (a) $\Lambda_{max}(\nu)$, for $\alpha=0$ (grey triangles) and $\alpha=\pi/2-0.1$ (black circles). (b) $\Lambda_{max}$ for a WS network as a function of the system size, $N$, and coupling strength, $d$, for rewiring probabilities (b) $p=0.1$, (c) $p=0.3$, and (d) $p=1.0$. The black dashed lines in (b)-(d) show the approximated critical curve, $\Lambda_{max}(\nu)=0$, obtained from the Wu-Chua conjecture (see Appendix C). Shades of red and blue correspond to a positive and negative sign of the MSF, respectively.
As seen earlier with chimera states, the system can still approach the synchronisation manifold in regions where it is unstable. In the following, we show that before transitioning from one dynamical regime to the other, the probability that the system exhibits anomalously high levels of synchronisation reaches maximum levels. We use this as means for characterizing the region of existence of chimera states and, thus epileptic-like events.
@@ II.3 Extreme events at the onset of phase transitions
By analysing the probability distribution (PDF) for the global order parameter, we studied the likelihood that an event of high synchronisation occurs spontaneously in the system. For this, we use the standard hydrodynamical criterion [37, 38, 39]: Given the PDF $P(A)$ of an observable $A$, a value $y$ of the observable is considered to be extreme if is larger than twice a threshold $A_{s}$, defined as the mean value of the highest tertile of $P(A)$, i.e., events with an abnormality index $I\equiv A/A_{s}>2$ (see Fig. 6a). To apply this criterion, we considered the observable $A(t)=-\log(1-R(t))$, which maps the domain of $R(t)$ from $[0,1]$ to $[0,+\infty)$. This ensures that the AI is always contained in the domain.
Throughout this section, we refer to $A$ as synchronisation amplitude, and the proportion of extreme events in a time series $A(t)$ is then defined as
@@
To illustrate how $p_{EE}$ changes with the coupling strength, we consider a system of $N=50$ oscillators in a WS network with $p=1$. As the coupling strength $d$ is increased (see Figs. 6b-c), the distribution $P(A)$ is altered significantly. The distribution transitions from a regime where extreme events of synchronisation amplitude are present (Figs. 6b and d), to a region where no extreme events are observed (see Fig. 6c).
Figure 6: Extreme events on networks of FHN oscillators. (a) Time-series of synchronisation amplitude $A(t)$, for $N=50$ oscillators, $d=0.04$ and $p=1.0$. The dotted line show the value $2A_{s}$ above which the values of $A(t)$ are considered extreme events (shown in red). (b)-(d) Complementary cumulative distributions of $A$, for (b) $d=0.04$, (c) $d=0.08$, and (d) $d=0.11$. The vertical dotted lines show the value $2A_{s}$.
A systematic analysis of the proportion of extreme events revealed a non-monotonic behaviour of $p_{EE}$ as the coupling strength is increased from zero up to the region of full synchronisation, as shown in Fig. 7a for $N=50$ and $p=1$. For low couplings ($d<0.05$) the $p_{EE}$ is always positive, showing a peak of events prior to the transition to a region where no extreme events are observed. As $d$ is increased further, a second peak in the $p_{EE}$ is observed (for $d\approx 0.11$), above which the $p_{EE}$ again falls to zero, and remains at zero for all larger values of $d$.
To understand this behaviour is worth looking at the nature of the incoherent behaviour for values of $d$ below the synchronisation transition (as given by the MSF analysis in Fig. 5a). For this, we computed the complete set of $2N$ Lyapunov exponents $\left\{\lambda_{i}\right\}_{i=1}^{2N}$ for the system with $N=50$ and $p=1$, for all values of $d$ [40, 41, 35]. The maximum lyapunov exponent $\lambda_{0}$ as a function of $d$ is shown in Fig. 7b. Moreover, we computed the Kaplan-Yorke dimension, corresponding to an upper bound to the dimension of the strange attractor, which is defined as [41, 35]
@@
where $j$ is the largest index such that $\sum_{i=1}^{j}\lambda_{i}\geq 0$. This analysis shows that, as expected, the complete region of incoherence corresponds to a chaotic region, due to the positivity of its largest Lyapunov exponent. In addition, we observe that the peaks in the $p_{EE}$ as $d$ is increased, precedes sharp decrease in the Kaplan-Yorke dimension, as has been observed for similar measures in other systems [39]. For the system with $N=50$ oscillators, the first region of positive $p_{EE}$ ($d<0.05$) correspond to high-dimensional chaos (see $D_{KY}$ in Fig. 7b), with the magnitude of the largest Lyapunov exponents increasing with $d$ before suffering a drastic reduction, just as the peak in the $p_{EE}$ is passed (see $\lambda_{0}$ in Fig. 7b). Interestingly, the peak in $p_{EE}$ marks the onset for the appearance of epileptic-like events (see Fig. 7a). In the intermediate region, where $p_{EE}=0$ ($0.10.05<d<$ 0.1), the chaotic attractor is significantly contracted and low-dimensional chaos is observed, i.e. there is only one positive Lyapunov exponent. There, the system exhibits coexistence of coherent and incoherent phases and, thus chimeras and epileptic-like events are observed throughout (see Fig. 7a). The second peak in the $p_{EE}$, coincides with the critical curve ($\Lambda_{max}=0$) predicted by the MSF analysis (see Fig. 5d). In this transition, both the largest Lyapunov exponents and the Kaplan-Yorke dimension are further reduced, with $D_{KY}$ decrease to zero, indicating the stabilization of the synchronisation manifold.
Figure 7: Chaos analysis of coupled FHN oscillators. (a) Left axis: Proportion of extreme events (black circles) for different values of coupling strength $d$. Right axis: Frequency of epileptic-like events (red triangles), and order parameter (grey diamonds) as in Fig. 4a. (b) Left axis: Maximum Lyapunov exponent (black circles), right axis: Kaplan-Yorke dimension (grey diamonds). All data shown were obtained from considering a system of $N=50$ FHN oscillators on WS networks with $p=1.0$.
Overall, comparing the time-averaged global order parameter, the frequency of epileptic-like events, the Kaplan-Yorke dimension, and the proportion of extreme events (Figs. 7a-b), we observe that changes in $D_{KY}$, as well as the disappearance and appearance of EEs, is directly correlated with the critical values of $d$ where the transitions occur.
As such, we computed the $p_{EE}$ in the complete $N-d$ parameter space (see Fig. 8). Overlaying to this the transition curve obtained from the MSF, we observe that peaks in the proportion of extreme events accurately predict the transition to the synchronised phase. Comparing with the frequency of epileptic-like events, we observe that their appearance is preceded by a peak in $p_{EE}$, and their frequency is maximum in the intermediate region where $p_{EE}=0$.
Figure 8: Extreme events on WS networks of FHN oscillators. Proportion of extreme events, $p_{EE}$, as a function of the system size ($N$) and coupling strength $d$, for $p=1.0$. Pale colour indicates regions where $p_{EE}=0$. The dashed line shows the critical curve where the MSF changes sign, obtained through the Wu-Chua conjecture.
@@ III Conclusions and discussion
In this study, we explored the dynamics of coupled FitzHugh-Nagumo oscillators on Watts-Strogatz small-world networks. Our results show that network randomness and coupling strength significantly influence the occurrence of epileptic-like states, which resemble the transient synchronisation observed in real neural networks during seizures.
@@
Interestingly, epileptic-like events occur only within an intermediate range of coupling strengths and become more frequent as the network becomes larger provided that the rewiring probability is large enough (as was shown in Fig. 3b-c). This opens questions about the compartmentalization in animal brains, and the mechanisms of wiring in healthy and altered conditions. In humans, for instance, anatomical inspection of the brain allows to identify of the order of 90 distinct regions [42, 43], with a inhomogeneous (small-world-like) structural connectivity (see Fig. 1b). According to our theoretical observations, this brain organisation promotes the coexistence between coherent and incoherent phases, yet it remains unclear what are the mechanisms that lead to this organisation.
@@
Evaluating the master stability function allowed us to find the transition curve to synchronisation in the $N-d$ parameter spaces for different values of the rewiring probability. This curve separated the phase space into a synchronised and a chaotic phase, as demonstrates by the Lyapunov exponent analysis. Within the chaotic phase, we distinguished two regions, which became apparent with when analysing the proportion of extreme events in the (rescaled) order parameter: A fully chaotic region and a region of coexistence between an incoherent and a coherent phase, where chimera states and epileptic-like events were observed. We have shown that the transitions from chaos to intermittency, and subsequently to synchronisation, are preceded by rise in the frequency of synchronisations events, a behaviour that correlates with significant changes in the dimensionality of the system’s strange attractor.
@@
We note that experimentally, the computation of the Lyapunov exponents is challenging. However, computing the proportion of extreme events $p_{EE}$ requires only a sufficiently long time series in order to construct the PDF of a relevant observable, which may provide useful information of the distinct dynamical phases of the system. This type of characterization of the phase space could open the door for the potential analysis of real EEG data or larger system sizes, where access to the Lyapunov spectrum may be limited, but long time series of the neural dynamic are more readily available.
@@
In conclusion, this work contributes to a deeper understanding of the interplay between network structure, coupling strength, and neural dynamics. The findings may have broader implications for understanding neural processes, offering potential applications in modelling the conditions that lead to pathological states such as epilepsy. Future work could extend this analysis to dynamical network topologies and explore the effects of additional biological factors, such as noise and time delays, to better model realistic neural systems.

### v2

@@ II.1 Emergence of ESR synchronization events
In the following, we study more generally the region of existence of chimeras (including both ESR events and persisting chimeras), and in particular, the transitions between the distinct dynamical regimes.
Figure 3: Emergence of ESR states. (a)-(c) Frequency of ESR events in $(N,d)$-parameter space, for (a) $p=0.1$, (b) $p=0.3$, and (c) $p=1.0$. (d) Shows the distribution of durations of the ESR events in log-linear and (inset) log-log scale, (e) shows the distribution of waiting times between ESR events. Markers in (c) indicate the parameters analyzed in (d) and (e). Black arrows in (d) and (c) show the net effect on the distributions of increasing either $d$ or $N$, while keeping other parameters fixed.
II.2 Characterizing the N-d-p phase space
@@
To characterize the transitions between dynamical states in the system, we compute the time-averaged global order parameter ($\bar{R}$) as a function of the coupling strength in WS networks of different sizes and $p=1$ (see Fig. 4a). We note that regardless of the network size, $\bar{R}$ grows monotonically with $d$, exhibiting two clear transitions: One at $d\sim 0.05$ and another at $d\sim 0.15$, after which the system reaches full synchronization. Notably, for all system sized, in the range $d\sim[0.05,0.15]$, where chimeras are expected (i.e. in the intermittent and partially synchronized regime) the averaged order parameter exhibits a plateau, where $\bar{R}$ is approximately independent of $d$. Increasing the rewiring probability causes a significant decrease in the coupling threshold at which the oscillators become fully synchronized (see Fig. 4b-c).
To understand the changes in the transition to synchronization, in the following we perform a linear stability analysis of the synchronized solution.
@@
Figure 4: Phase transitions of coupled FHN oscillators. (a) Time-averaged order parameter, $\bar{R}$, as a function of the coupling strength, $d$, for WS networks with rewiring probability $p=1$ and varied sizes, $N$. Inset shows a close-up of the transition from intermittency to global synchronization. (b)-(d) $\bar{R}$ in the $(N,d)$-parameter space, for (b) $p=0.1$, (c) $p=0.3$, and (d) $p=1.0$.
@@ II.2.1 Linear stability analysis of the synchronous solution
A linear stability analysis for the synchronized solution can be achieved by evaluating the master stability function (MSF) of Eq. (1) [46, 3, 47] (see derivation in Appendix B). The MSF is defined as the maximum Lyapunov exponent $\Lambda_{max}(\nu)$ associated with the linearized dynamical system describing the time evolution of the perturbations transverse to the synchronous manifold $\mathcal{S}$. The MSF defines the stability of $\mathcal{S}$ as a function of a single parameter $\nu=\gamma_{2}d$, where $d$ is the coupling strength and $\gamma_{2}$ is the second smallest eigenvalue of the network Laplacian matrix, known as algebraic connectivity [48]. When $\Lambda_{max}(\nu)>0$, $\mathcal{S}$ is unstable, while $\Lambda_{max}(\nu)<0$ provides a necessary condition for the stability of $\mathcal{S}$ [3]. In Fig. 5a, we show that $\Lambda_{max}(\nu)$ is positive for a range of values of $\nu$ for $\alpha=\pi/2-0.1$, however, when $\alpha=0$ the MSF is always negative and neither chimeras nor chaotic behavior can be observed.
Consistent with the observations of the time-averaged global order parameter (Fig. 4b-d), lowering the rewiring probability $p$ causes a shift in the critical curve $\Lambda_{max}(\nu)=0$ towards larger values of $d$ (see Fig. 5b-d). We note that the critical curve was accurately predicted by the Wu-Chua conjecture (see Appendix C), delimiting the values above which the synchronized manifold is stable (see dashed line in Fig. 5b-d).
The MSF analysis shows that the change in the transition line to synchronization comes from changes in the topological features of the network, which enter in the MSF through the second smallest eigenvalue of the Laplacian matrix of the connectivity network. With this, we have characterized the transition from the incoherent to the coherent regime. However, the transition from chaos to intermittency is not captured by the MSF approach.
Figure 5: MSF of FHN coupled oscillators. (a) $\Lambda_{max}(\nu)$, for $\alpha=0$ (gray triangles) and $\alpha=\pi/2-0.1$ (black circles). (b) $\Lambda_{max}$ for a WS network as a function of the system size, $N$, and coupling strength, $d$, for rewiring probabilities (b) $p=0.1$, (c) $p=0.3$, and (d) $p=1.0$. The black dashed lines in (b)-(d) show the approximated critical curve, $\Lambda_{max}(\nu)=0$, obtained from the Wu-Chua conjecture (see Appendix C). Shades of red and blue correspond to a positive and negative sign of the MSF, respectively.
As seen earlier with chimera states, the system can still approach the synchronization manifold in regions where the synchronized solution is unstable. In the following, we show that before transitioning from one dynamical regime to another, the probability that the system exhibits anomalously high levels of synchronization reaches maximum levels. We use this as a means for characterizing the region of existence of chimera states and, thus, ESR events.
@@ II.3 Extreme events at the onset of phase transitions
By constructing the probability distribution (PDF) of the global order parameter, we studied the likelihood that extreme events of anomalously high synchronization occur as a function of the system parameters. For this, we use the standard hydrodynamical criterion [49, 50, 51]: Given the PDF $P(A)$ of an observable $A$, a value $y$ of the observable is considered to be extreme if it is larger than $2A_{s}$, with $A_{s}$ defined as the mean of the third tertile of $P(A)$, i.e., events with an abnormality index $AI\equiv A/A_{s}>2$ (see Fig. 6a). To apply this criterion, we considered the observable $A(t)=-\log(1-R(t))$, which maps the domain of $R(t)$ from $[0,1]$ to $[0,+\infty)$. This ensures that the $AI$ is always contained in the domain. Throughout this section, we refer to $A$ as synchronization amplitude. The proportion of EEs for a time series $A(t)$ is defined as
@@
To illustrate how $p_{EE}$ changes with the coupling strength, we consider a system of $N=50$ oscillators in a WS network with $p=1$. As the coupling strength $d$ is increased from $d=0.04$ to $d=11$, the system transitions from chaotic ($d=0.04$), to intermittent ($d=0.08$), to partial synchronization ($d=0.11$), and the distribution $P(A)$ changes significantly (see Figs. 6b-c). In the chaotic regime and in the transition to synchronization, the system exhibits a positive proportion of extreme events (see red markers in Figs. 6b and 6d). However, in the intermediate region, where ESR events are observed, the proportion of EEs is zero (see Fig. 6c).
Figure 6: EEs on networks of FHN oscillators. (a) Time-series of synchronization amplitude $A(t)$, for $N=50$ oscillators, $d=0.04$ and $p=1.0$. The dotted line show the value $2A_{s}$ above which the values of $A(t)$ are considered EEs (shown in red). (b)-(d) Complementary cumulative distributions of $A$, for (b) $d=0.04$, (c) $d=0.08$, and (d) $d=0.11$. The vertical dotted lines show the value $2A_{s}$.
A systematic analysis of the proportion of EEs revealed a non-monotonic behavior of $p_{EE}$ as the coupling strength is increased from zero up to the region of full synchronization, as shown in Fig. 7a for $N=50$ and $p=1$. For low couplings ($d<0.05$) the $p_{EE}$ is always positive, showing a peak of events prior to the transition to intermittency, where no EEs are observed. As $d$ is increased further, a second peak in the $p_{EE}$ is observed (for $d\approx 0.11$), above which the system reaches full synchronization and $p_{EE}$ again falls to zero.
To understand this behavior, we looked in more detail at the dynamic before the synchronization transition. For this, we computed the complete set of $2N$ Lyapunov exponents $\left\{\lambda_{i}\right\}_{i=1}^{2N}$ for the system with $N=50$ and $p=1$, for all values of $d$ [52, 53, 47]. The maximum Lyapunov exponent $\lambda_{0}$ as a function of $d$ is shown in Fig. 7b, which remain positive below the synchronization transition. We also computed the Kaplan-Yorke dimension, corresponding to an upper bound to the dimension of the strange attractor, which is defined as [53, 47]
@@
where $j$ is the largest index such that $\sum_{i=1}^{j}\lambda_{i}\geq 0$. For a system with $N=50$ oscillators, the first region of positive $p_{EE}$ ($d<0.05$) corresponds to high-dimensional chaos (see $D_{KY}$ in Fig. 7b), with the magnitude of the largest Lyapunov exponent increasing with $d$ before suffering a drastic reduction, as the peak in the $p_{EE}$ is passed (see $\lambda_{0}$ in Fig. 7b). The peak in $p_{EE}$ is followed by a reduction of the Kaplan-Yorke dimension and marks the onset of intermittency, where ESR events are observed (see Fig. 7a). Both the intermittent and the partially synchronized regime are characterized by low-dimensional chaos (with only one positive Lyapunov exponent). The second peak in the $p_{EE}$ coincides with the critical curve ($\Lambda_{max}=0$) predicted by the MSF analysis (see Fig. 5d), after which the largest Lyapunov exponent and the Kaplan-Yorke dimension are further reduced.
Figure 7: EEs and Lyapunov analysis. (a) Proportion of EEs, frequency of ESR events (normalized by its maximum value) and order parameter (reproduced from Fig. 4a), as a function of the coupling strength $d$. (b) Maximum Lyapunov exponent and Kaplan-Yorke dimension as a function of the coupling strength $d$. All data shown were obtained from considering a system of $N=50$ FHN oscillators on WS networks with $p=1.0$. (a-b) The shaded region indicates the existence of chimeras, the transition point from the region of chaos to chimeras is identified by the appearance of ESR events, while the transition point from chimeras to synchronization corresponds to $d=0.116$, as predicted by the MSF analysis.
We then computed the $p_{EE}$ in the complete $N-d$ parameter space (see Fig. 8). Here, we notice that in the chaotic regime $p_{EE}$ is always positive. As noted before, at the onset of intermittency, the proportion of EEs exhibits a peak, decaying to zero once the intermittent region is reached. By further increasing the coupling strength, a second peak is observed in the proportion of EEs, accurately matching the transition curve obtained from the MSF analysis for the stabilization of the synchronization manifold. We note that our analysis based on the proportion of EEs does not distinguish between the intermittent regime (i.e. coexistence of chaos and chimeras) and that of partial synchronization (in which chimeras persist for the whole duration of the simulation).
@@ III Conclusions and discussion
In this study, we explored the dynamics of coupled FitzHugh-Nagumo oscillators on Watts-Strogatz small-world networks. Our results show that the network randomness and coupling strength significantly influence the occurrence of ESR states, which resemble the transient synchronization observed in biological neural networks during absence epileptic seizures. Interestingly, ESR events occur only within an intermediate range of coupling strengths and become more frequent and of shorter duration as the network becomes larger, provided that the rewiring probability is large enough (as shown in Fig. 3b-c). This opens questions about the compartmentalization in animal brains and the mechanisms of wiring in healthy and altered conditions. In humans, for instance, anatomical inspection of the brain enables the identification of approximately 90 distinct regions [54, 55]. According to our theoretical observations, this brain organization promotes the coexistence of coherent and incoherent phases, however, it remains unclear which are the mechanisms that lead to this specific organization.
Figure 8: EEs on WS networks of FHN oscillators. Proportion of EEs ($p_{EE}$) as a function of the system size $N$ and coupling strength $d$, for $p=1.0$. Pale color indicates regions where $p_{EE}=0$. The dashed line shows the critical curve where the MSF changes sign, obtained through the Wu-Chua conjecture.
@@
Evaluating the master stability function allowed us to find the transition curve to synchronization in the $N-d$ parameter space for different values of the rewiring probability $p$. This curve separated the phase space into a synchronized and a chaotic phase, as demonstrated by the Lyapunov exponent analysis. Within the chaotic phase, we distinguished two regions, which became apparent when analyzing the proportion of EEs in the (rescaled) order parameter: A fully chaotic region and a region of coexistence between an incoherent and a coherent phase, where both chimera states and ESR events were observed. We have shown that the transitions from chaos to intermittency, and subsequently to synchronization, are preceded by a peak in the frequency of extreme synchronization events, a behavior that correlates with significant changes in the dimensionality of the system’s strange attractor.
@@
In general, computing the proportion of EEs $p_{EE}$ requires only a sufficiently long time series in order to construct the PDF of a relevant observable, which may provide useful information of the distinct dynamical phases of the system. This type of characterization of the phase space could open the door for the potential analysis of real EEG data, where time series of neural dynamics are more readily available.
@@
In conclusion, this work contributes to a deeper understanding of the interplay between network structure, coupling strength, and neural dynamics. The findings may have broader implications for understanding neural processes, offering potential applications in modeling the conditions that lead to pathological states such as ESR intermittency and synchronous states. Future work could extend this analysis to dynamical network topologies and explore the effects of additional biological factors such as noise and time delays, to better model realistic neural systems.